# Question-Driven Analysis and Synthesis: Building Interpretable Thematic Trees with LLMs for Text Clustering and Controllable Generation

### v1

@@ Abstract
Unsupervised analysis of text corpora is challenging, especially in data-scarce domains where traditional topic models struggle. While these models offer a solution, they typically describe clusters with lists of keywords that require significant manual effort to interpret and often lack semantic coherence. To address this critical interpretability gap, we introduce Recursive Thematic Partitioning (RTP), a novel framework that leverages Large Language Models (LLMs) to interactively build a binary tree. Each node in the tree is a natural language question that semantically partitions the data, resulting in a fully interpretable taxonomy where the logic of each cluster is explicit. Our experiments demonstrate that RTP’s question-driven hierarchy is more interpretable than the keyword-based topics from a strong baseline like BERTopic. Furthermore, we establish the quantitative utility of these clusters by showing they serve as powerful features in downstream classification tasks, particularly when the data’s underlying themes correlate with the task labels. RTP introduces a new paradigm for data exploration, shifting the focus from statistical pattern discovery to knowledge-driven thematic analysis. Furthermore, we demonstrate that the thematic paths from the RTP tree can serve as structured, controllable prompts for generative models. This transforms our analytical framework into a powerful tool for synthesis, enabling the consistent imitation of specific characteristics discovered in the source corpus.
@@ Introduction
    By comparing our method with a BERTopic baseline, we show that the thematic hierarchies generated by RTP are easier to interpret and more often linked to actionable information within the dataset.
@@
    We establish the quantitative utility of the generated clusters by showing they serve as effective features in downstream classification tasks.
@@
    We introduce and validate a novel application: using the paths of the tree as a structured “thematic signature” for Controllable Thematic Generation (CTG). Our results reveal that this technique provides unparalleled control over the semantic content of generated text, effectively solving a key challenge in controllable synthesis.
@@
    Global Corpus Sampling. First, we draw a single, global random sample from the entire corpus. This smaller, sampled dataset is then used as the input for the complete recursive partitioning process. The sample size is a key hyperparameter governing the trade-off between representativeness and computational cost.
@@
    Document Summarization. Second, if the documents within our global sample are themselves very long, we can perform a one-time summarization step. Each document in the sample is summarized using an LLM call. We treat the use of summarization as a boolean hyperparameter.
@@
At each node of the tree, starting with the root node containing the entire sampled dataset, we prompt an LLM to generate a single, divisive thematic question. The prompt contains a representative subset of documents from the current partition and instructs the LLM to create a yes/no question that best splits these documents into two coherent, roughly equal-sized groups.
@@
Once a question is generated, we prompt the LLM to generate the answer to the question for each document within the partition. To mitigate stochastic LLM errors, we employ a majority voting mechanism:
@@
The number of votes, N, is a hyperparameter allowing a trade-off between robustness and token count. Documents are then split into a “yes” child node and a “no” child node based on their final assigned answer.
@@
Both D and K are hyperparameters of the framework.
@@
Controllable Thematic Generation (CTG) leverages the explicit structure of the RTP tree to generate texts that adhere to specific, user-defined themes. The process begins by selecting a target leaf node that represents the desired thematic content. We then extract the “thematic signature” of this node, defined as the unique sequence of questions and answers along the path from the root. This sequence is then translated into a set of structured constraints that are incorporated into a prompt to guide a generative LLM.
@@
Qualitative Analysis of Thematic Trees
@@
We generated thematic trees for each dataset to perform a qualitative analysis of RTP’s behavior [^4] Parameters: sample size 100, $N=4$, $D=5$, $K=3$, no summarization. They are shown in figures 2 (IMDB), 3 (Yelp Reviews), and 4 (AG-News). We immediately note that the questions seem to be related to the dataset content.
[root] Does the text primarily provide a positive review of the film or work discussed?
      [yes] Does the text describe the production or plot of a movie, show, or documentary that involves or depicts a romantic relationship between characters?
      [yes] Does this text describe a film where the main character faces a personal moral or ethical struggle?
      [no] Does the text describe a romantic relationship as a central theme in the movie?
      [yes] Does the text mention a prominent female character in the story?
      [no] Does the text discuss the cinematography or visual presentation of the film or show?
      [yes] Does the text mention a documentary or a film about saving the planet or environmental issues?
      [no] Does the text describe the film in a very positive light overall?
      [no] Does the text describe a film or show as being emotionally impactful?
      [yes] Does the text recommend watching the movie or show?
      [no] Does the text mention a television show or series?
      [no] Does the text describe the movie as disappointing or poorly executed?
      [yes] Does the reviewer criticize the script for being poorly written or uninteresting?
      [yes] Does the text have a positive overall assessment of the movie?
      [no] Does the text have a positive overall assessment of the movie?
      [no] Does the text mention disappointment with the film’s acting or storyline?
      [yes] Does the text criticize the film’s lack of coherence or clear direction?
      [no] Does the text describe the movie as boring?
      [no] Does the text mention a train?
      [no] Does the text mention a specific film director by name?
      [no] Does the text critique an unjust or exaggerated portrayal of characters or events?
Figure 2: Thematic tree generated for the IMDB dataset.
[root] Does the text mention the author having a positive experience with the food or services provided?
      [yes] Does the review mention being satisfied with the food at the establishment?
      [yes] Does the text mention a specific dish or food item that the author enjoyed at the establishment?
      [yes] Does the text mention a specific dish or menu item that the author enjoyed?
      [yes] Does the text mention a specific dish or menu item that the author enjoyed?
      [no] Does the text mention positive aspects of the customer service or staff?
      [yes] Does the reviewer mention exceptional or notable service in the text?
      [no] Does the text mention a specific issue or drawback related to the described experience or place?
      [no] Does the text mention dissatisfaction with food or drink quality?
      [yes] Does the review mention an issue with the food’s preparation or quality?
      [yes] Does the review mention an issue with the food’s preparation or quality?
      [no] Does the text describe a place where a purchase can be made?
      [yes] Does the text mention a specific price for a product or service?
      [no] Does the text mention experiencing poor customer service?
      [yes] Does the text describe an experience with poor service?
      [yes] Does the text describe an experience with poor service?
      [yes] Does the text describe an experience with poor service?
      [no] Does the text describe a positive experience with the food at the establishment?
      [no] Is this review primarily focused on dissatisfaction with the food or beverage quality?
      [yes] Does the text describe significant dissatisfaction with the food quality?
      [no] Does the text describe an experience involving food or drink?
Figure 3: Thematic tree generated for the Yelp Reviews dataset.
[root] Does the text describe a significant organizational, economic, or structural change?
      [yes] Does the text mention a plan or agreement to cut jobs, operations, or resources?
      [no] Does the text mention financial losses or challenges faced by a company?
      [yes] Does the text mention financial troubles or losses for the company involved?
      [yes] Does the text report a resolution or agreement in response to a financial or operational challenge?
      [no] Does the text mention a company cutting, reducing, or altering any of its offerings or practices?
      [yes] Does the text mention a reduction or change being made to a product or practice?
      [no] Does the text mention a merger or collaboration between two entities?
      [no] Does the text mention sports or athletes?
      [yes] Does the text mention a victory or winning event?
      [yes] Does the text mention a specific individual achieving victory or recognition?
      [yes] Does the text mention a victory or win in a sports competition?
      [no] Does the text describe a decisive event or moment in a sports context?
      [yes] Does the text specifically mention a score or ranking?
      [no] Does the text mention an injury?
      [no] Does the text explicitly mention a financial market or economic activity?
      [yes] Does the text mention oil prices or crude oil?
      [yes] Does the text mention oil prices changing?
      [no] Does the text mention a financial or economic statistic or figure?
      [no] Does the text mention a government action or decision?
      [yes] Does the text mention an instance of a government or official body negotiating or responding to a conflict or dispute?
      [no] Does the text mention a specific event occurring in nature or the environment?
Figure 4: Thematic tree generated for the AG-News dataset.
A qualitative examination of the trees generated by RTP reveals several key characteristics of our method, highlighting both its strengths in thematic discovery and its interesting emergent behaviors.
@@
Thematic Redundancy. We also noted instances of similar or repeated questions appearing in the tree, either in parent-child relationships or between sibling nodes. This is a natural consequence of the algorithm’s design, where question generation for each node is an independent process that only considers the local document partition. The model has no memory of the questions asked previously in the hierarchy. While this can lead to redundancy, it also demonstrates the robustness of the LLM in identifying the most salient thematic split at different levels of granularity.
@@
To provide a fair comparison with a standard topic model, we executed BERTopic in the same dataset sample used in the experiments with RTP. The results are shown in Tables 1, 2, and 3. Following BERTopic convention, topics are numbered from 0 to the number of topics found, and topic 1-1 represents outliers.
@@
Analysis
Text Classification
Setup
To obtain a baseline result, we fine-tune a pre-trained DistilBERT model (Sanh et al. 2019) on the exact same small data sample used to generate the RTP tree. This ensures a fair comparison, isolating the benefit of RTP’s structure against the fine-tuning of a powerful neural network on the same limited data. For context, we also report the state-of-the-art (SoTA) performance from models on huggingface.com  [^5] IMDB: https://huggingface.co/DeepAxion/distilbert-imdb-sentiment; Yelp Reviews: https://huggingface.co/BexRedpill/distilbert-on-yelp-reviews-full-epoch-2; AG-News: https://huggingface.co/JiaqiLee/bert-agnews, which are typically trained on the full dataset.
Results
Our primary finding is that RTP significantly outperforms the fine-tuned DistilBERT baseline in IMDB, and surpasses the performance of a model trained on the full dataset ($p<0.05$, one-sample t-test). The performances of RTP and the baseline are similar in Yelp Reviews. However, on the topic-based AG News dataset, the fine-tuned DistilBERT baseline is superior. This brings forward a key characteristic of our method.
Model IMDB Yelp Reviews AG-News
SoTA (reference) .94 .65 .94
DistillBERT (baseline) .81(.03) .39(.07) 0.87(.01)
RTP (ours) 96(.02) .40(.12) .64(.04)
RTP-T (ours) .96(.02) .39(.04) .52(.10)
RTP-E (ours) .91(.03) .37(.08) .61(.06)
RTP-TE (ours) .89(.04) .41(.05)* .51(.07)
Table 4: Accuracies for each model. SoTA is the highest accuracy reported in models found in huggingface.com. The RTP classifiers are identified on whether they use summarization on training (T) or on evaluation (E). The highest mean accuracy (excluding SoTA) is marked in bold. Values are $\mu(2\sigma)$. Asterisks denote non-significant differences ($p>.05$) compared to baseline.
Analysis
Controllable Thematic Generation
Setup
Results
Strategy IMDB Yelp Reviews AG-News
Uncontrolled .37 .51 -.01
Few-shot (baseline) .45 .55 .86
CTG (ours) .41 .52 .34
Strategy IMDB Yelp Reviews AG-News
Uncontrolled .07 .00 .00
Few-shot (baseline) .04 .03 .08
CTG (ours) .60 .39 .31
Table 6: Node accuracy evaluation of each generation strategy.
Analysis
Conclusion

### v2

@@ Abstract
Unsupervised analysis of text corpora is challenging, especially in data-scarce domains where traditional topic models struggle. While these models offer a solution, they typically describe clusters with lists of keywords that require significant manual effort to interpret and often lack semantic coherence. To address this critical interpretability gap, we introduce Recursive Thematic Partitioning (RTP), a novel framework that leverages Large Language Models (LLMs) to interactively build a binary tree. Each node in the tree is a natural language question that semantically partitions the data, resulting in a fully interpretable taxonomy where the logic of each cluster is explicit. Our experiments suggest that RTP’s question-driven hierarchy is more interpretable than the keyword-based topics from a baseline like BERTopic. Furthermore, we show that formed clusters can serve as powerful features in downstream classification tasks, particularly when the data’s underlying themes correlate with the task labels. RTP introduces a new paradigm for data exploration, shifting the focus from statistical pattern discovery to knowledge-driven thematic analysis. Furthermore, we demonstrate that the thematic paths from the RTP tree can serve as structured, controllable prompts for generative models. This transforms our analytical framework into a powerful tool for synthesis, enabling the consistent imitation of specific characteristics discovered in the source corpus.
@@ Introduction
    We demonstrate that RTP generates thematic hierarchies based on explicit, natural language questions. By contrasting our method with a keyword-based baseline like BERTopic, we show that this question-driven approach offers a transparent, rule-based form of interpretability where the logic of each partition is human-readable by design.
@@
    We investigate the quantitative utility of the generated clusters as features for downstream classification. Our preliminary analysis highlights their potential in settings where the discovered themes align with the task’s underlying labels.
@@
    We introduce a novel application for the resulting thematic paths: Controllable Thematic Generation (CTG). We demonstrate how these “thematic signatures” can be used as structured prompts to exert fine-grained, rule-based control over the semantic content of generated text, distinguishing this approach from style-imitation techniques.
@@
    Global Corpus Sampling. First, we draw a single, global random sample from the entire corpus. This smaller, sampled dataset is then used as the input for the complete recursive partitioning process. The sample size $S$ is a key hyperparameter governing the trade-off between representativeness and computational cost. Also,
@@
    Document Summarization. Second, if the documents within our global sample are themselves very long, we can perform a one-time summarization step. Each document in the sample is summarized using an LLM call, using the prompt in Figure 2 as system instructions and the text to be summarized as user instructions. We treat the use of summarization as a boolean hyperparameter.
System Prompt Summarize the following text into summary with a maximum of {max_words} words. The summary should be concise and capture the main points of the text. Reply solely with the summary, without any additional text or formatting. Your answer should be a JSON object following the schema: {schema_str} User Prompt Text: {text}
Figure 2: Prompts used for summarization. The text data is provided to the LLM as a user prompt. By changing the max_words parameter, we can control the summary length. The JSON schema used for prompting helps parsing the answer.
@@
At each node of the tree, starting with the root node containing the entire sampled dataset, we prompt an LLM to generate a single, divisive thematic question, using the system instructions shown in Figure 3. The user prompt contains a representative subset of documents from the current partition and instructs the LLM to create a yes/no question that best splits these documents into two coherent, roughly equal-sized groups.
System Prompt You are an analyst. You will receive a list of texts formated as a list. You will propose one yes/no question whose answer can be determined from each individual text. The question must be answerable “yes” to half of these texts, and “no” to the other half. The question must be concise and clear, allowing for a straightforward yes/no answer for each text. Do not use any external knowledge or assumptions. Your answer must be determined only from the texts provided. You must reply following a json in the schema: {schema_str} User Prompt Texts: {texts}
Figure 3: Prompts used for question generation. The text data is provided to the LLM as a user prompt containing a Python-style list of texts. The JSON schema used for prompting helps parsing the answer.
@@
Once a question is generated, we prompt the LLM to generate the answer to the question for each document within the partition, using the prompt shown in Figure 4.
System Prompt You will receive a text and a question about it. Answer the question as yes (true) or no (false) based solely on the text. Do NOT use any external knowledge or assumptions. Your answer must be determined only from the text provided. If the question is not answerable from the text, respond with ”false”. Ignore assumptions or implications not explicitly stated in the text. Reply with a JSON object following the schema: {schema_str} User Prompt Text: {text} Question: {question}
Figure 4: Prompt used for data partitioning. The text data is provided to the LLM as a user prompt containing a Python-style list of texts. The JSON schema used for prompting helps parsing the answer.
To mitigate stochastic LLM errors, we employ a majority voting mechanism:
@@
The number of votes, N, is a hyperparameter allowing a trade-off between robustness and token count. This parameter is more relevant when the questions and related documents are more ambiguous (for example: asking whether a news article narrating a football game is about sports is straightforward, whereas asking whether that same news article narrates overall positive news can be different depending on the reader’s opinion about the winning team). Documents are then split into a “yes” child node and a “no” child node based on their final assigned answer.
Draws are arbitrarily assumed to be a “yes”. This could lead to biasing towards positive answers. However, draws tend to only occur if questions are ambiguous in relation to the analyzed text, which means either answer would be equally suitable.
@@
Both $D$ and $K$ are hyperparameters of the framework. Because the main content of the prompts is composed of texts from the dataset, and these texts are partitioned in each split, we can assume that the computational cost, in tokens, for this recursion is $O(SD)$, where $S$ is the sample size and $D$ is the maximum tree depth. $K$ can be chosen as a regularization parameter avoiding recursing into branches with too few elements, similarly to the regularization of decision trees.
Importantly, this recursive splitting technique means that, if the trees are balanced, then the maximum depth of the tree is $M_{d}=\lfloor\log_{2}S\rfloor+1$. This is an important feature of this technique, as it means that a larger sample allows asking more questions about the same
@@
Controllable Thematic Generation (CTG) leverages the explicit structure of the RTP tree to generate texts that adhere to specific, user-defined themes. The process begins by selecting a target leaf node that represents the desired thematic content. We then extract the “thematic signature” of this node, defined as the unique sequence of questions and answers along the path from the root. This sequence is then translated into a set of structured constraints that are incorporated into a prompt to guide a generative LLM, leading to a prompt as shown in Figure 5.
System Prompt You are a skillful writer. You are writing under the following context: [context describing the dataset idea, e.g., this is a movie review] Ensure the material is coherent and relevant to the questions asked. You will be asked to generate material based on questions and answers Reply with a json string that containing the generated material in the following schema: {schema_str} User Prompt Questions and answers: [list of questions and answers leading to the specific leaf node used for generation]
Figure 5: Prompt used for CTG. The JSON schema used for prompting helps parsing the answer.
@@
Hyperparameters
In these experiments, we used the following hyperparameters:
  • Sample size $S=100$, which we found to be high enough for a reasonable balanced tree depth $M_{d}=7$, and, simultaneously, allowing all elements withing the sample to fit within the LLM’s context window;
  • Maximum depth $D=5$, which is less than $M_{d}$ and aims to produce nodes with an average of $4$ samples;
  • Minimum elements per node $K=3$, avoiding leaf nodes that are too specific and become non-representative.
  • Number of answers $N=4$, aiming to stabilize answers without significantly increasing the token count.
These hyperparameters were chosen to favor particular sizes for the trees. A full empirical study of their effect in the tree induction process is a key direction for future work.
Analysis of Thematic Trees
We generated thematic trees for each dataset to perform a qualitative analysis of RTP’s behavior. The token usage is shown in Table 1. As we can see, token usage for the experiments was in the range of $10^{5}$ to $10^{6}$, hence costing less than 1 USD with current API princing.
Dataset Q A Total
IMDB 135204 231036 366240
Yelp 94629 192993 287622
AGNews 29615 122667 152282
Table 1: Token usage for each experiment in the question creation stage (Q), the question answering stage (A), and total.
@@
They are shown in figures 6 (IMDB), 7 (Yelp Reviews), and 8 (AG-News) (located in the appendix). We immediately note that the questions seem to be related to the dataset content.
Alignment with ground truth labels
To quantitatively analyze the alignment between RTP’s unsupervised partitions and the ground-truth labels, we employ two complementary metrics for each node: Information Entropy and Node Purity.
First, we use Information Entropy, a standard measure of impurity from information theory. A low entropy value indicates that a node contains documents from a limited number of classes, with a value of 0 representing a perfectly ”pure” node. It is calculated as:
$H(S)=-\sum_{i=1}^{c}p_{i}\log_{2}(p_{i})$ (1)
where $S$ is the set of documents in a given node, $c$ is the number of unique classes, and $p_{i}$ is the proportion of documents in $S$ belonging to class $c_{i}$.
Second, to provide a more direct and intuitive measure, we calculate Node Purity as the proportion of the majority class within the node:
$\text{Purity}(S)=\max_{i}(p_{i})$ (2)
While entropy captures the overall class distribution, purity offers a straightforward assessment of the dominance of a single class. Together, these metrics allow us to quantify the emergent correlation between the semantic themes discovered by the LLM and the latent class structure of the data. For instance, a high purity and low entropy in a leaf node indicate that the path of questions leading to it has successfully isolated a thematically coherent group that also shares a class label.
Since RTP is an unsupervised framework that receives no label information during tree construction, perfect purity in all leaves is not the expected outcome. Instead, these metrics provide a quantitative lens that complements the following qualitative examination of the trees, which reveals the specific nature of these thematic discoveries and other emergent behaviors of our method.
@@
Thematic Redundancy. We also noted instances of similar or repeated questions appearing in the tree, either in parent-child relationships or between sibling nodes. The redundant questions seem to appear in nodes whose purity is higher, which indicates this is linked to biases found within the nodes themselves. This is a natural consequence of the algorithm’s design, where question generation for each node is an independent process that only considers the local document partition. The model has no memory of the questions asked previously in the hierarchy. While this can lead to redundancy, it also demonstrates the robustness of the LLM in identifying the most salient thematic split at different levels of granularity.
@@
To provide a fair comparison with a standard topic model, we executed BERTopic in the same dataset sample used in the experiments with RTP. The results are shown in Tables 5, 6, and 7, provided in the Appendix. Following BERTopic convention, topics are numbered from 0 to the number of topics found, and topic 1-1 represents outliers.
Analysis
Text Classification
Setup
To obtain a baseline result, we fine-tune a pre-trained DistilBERT model (Sanh et al. 2019) on the exact same small data sample used to generate the RTP tree. This ensures a fair comparison, isolating the benefit of RTP’s structure against the fine-tuning of a powerful neural network on the same limited data. For context, we also report the state-of-the-art (SoTA) performance among all models available on huggingface.com  [^4] We manually sought models fine-tuned in the target datasets IMDB: https://huggingface.co/DeepAxion/distilbert-imdb-sentiment; Yelp Reviews: https://huggingface.co/BexRedpill/distilbert-on-yelp-reviews-full-epoch-2; AG-News: https://huggingface.co/JiaqiLee/bert-agnews. These results are compatible with those found in recent Kaggle competition leaderboards like https://www.kaggle.com/competitions/sentiment-analysis-on-imdb-dataset-2025/leaderboard and https://www.kaggle.com/competitions/enap-bert-para-classificao/leaderboard., which are typically BERT-based transformers (distillBERT for IMDB and Yelp, BERT for AG-News) trained on the full dataset. We include these scores not as a direct competitive benchmark, as our protocol is intentionally data-scarce, but as a reference point to contextualize our results.
Results
Our primary finding is that RTP significantly outperforms the fine-tuned DistilBERT baseline in IMDB. The performances of RTP and the baseline are similar in Yelp Reviews. However, on the topic-based AG News dataset, the fine-tuned DistilBERT baseline is superior.
Model IMDB Yelp Reviews AG-News
SoTA (reference) .94 .65 .94
DistillBERT (baseline) .81(.03) .39(.07) .87(.01)
RTP (ours) 96(.02) .40(.12) .64(.04)
RTP-T (ours) .96(.02) .39(.04) .52(.10)
RTP-E (ours) .91(.03) .37(.08) .61(.06)
RTP-TE (ours) .89(.04) .41(.05)* .51(.07)
Table 2: Accuracies for each model. SoTA is the highest accuracy reported in models found in huggingface.com. The RTP classifiers are identified on whether they use summarization on training (T) or on evaluation (E). The highest mean accuracy (excluding SoTA) is marked in bold. Values are $\mu(c)$, where $c$ is the $95\%$ confidence interval. Asterisks denote non-significant differences ($p>.05$ using a t-test) compared to baseline.
Analysis
Controllable Thematic Generation
Setup
Results
Strategy IMDB Yelp Reviews AG-News
Uncontrolled .37 .51 -.01
Few-shot (baseline) .45 .55 .86
CTG (ours) .41 .52 .34
Strategy IMDB Yelp Reviews AG-News
Uncontrolled .07 .00 .00
Few-shot (baseline) .04 .03 .08
CTG (ours) .60 .39 .31
Table 4: Node accuracy evaluation of each generation strategy.
Analysis
Conclusion
Appendix A Appendix: Thematic trees for all executions
Figure 6: Thematic tree generated for the IMDB dataset.
Figure 7: Thematic tree generated for the Yelp Reviews dataset.
Figure 8: Thematic tree generated for the AG-News dataset.
Appendix B Appendix: Topics generated by BERTopic